You toss 3 coins. What is the probability that you get exactly 2 heads, given that you get at least one head?
Answer by stanbon(48568)

You toss 3 coins. What is the probability that you get exactly 2 heads, given that you get at least one head?
-----------------------------
Formula: P(A|B) =[P(A and B)] / P(B)
P(2 heads on 3 tosses | get at least one head)
--------------
P(at least one head) = 1-P(no heads in three) = 1-1/8 = 7/8
------------------------
P(2 heads AND at least one head) = P(3 heads) = 1/8
--------------------------
So< P(2 heads on 3 toses | get at least one head) = [1/8]/[7/8] 1/7
=====================
Cheers,
Stan H.